OpenStudy (anonymous):

the ratio of red marbles to blue marbles in a bag is 4:7. there ARE 16 RED MARBLES HOW MANY BLUE MARBLES ARE THERE

OpenStudy (texaschic101):

what you got here is a proportion first let me explain that the ratio 4:7 can also be written as 4/7. 4/7 = 16/x ---(4 red to 7 blue = 16 red to x blue) as with any proportion, we now cross multiply (4)(x) = (16)(7) 4x = 112 -- now divide both sides by 4 x = 28 we can even check to see if our answer is correct.. 4/7 = 16/28 -- cross multiply (4)(28) = (7)(16) 112 = 112 (correct) so if he has 16 red marbles, then he has 28 blue ones any questions ?

TheSmartOne (thesmartone):